OpenStudy (anonymous):

Find the x- and y- intercepts of the line that passes through the given points. (-5,-5) and (5,-1)

OpenStudy (danjs):

calculate the slope first

OpenStudy (anonymous):

@DanJS The slope is 2/5 I think?

OpenStudy (danjs):

correct

OpenStudy (danjs):

now use that slope, and one of the points in the point slope forumula

OpenStudy (danjs):

y - y1 = m(x-x1)

OpenStudy (anonymous):

\[ \frac{-5-(y_1)}{-5-(0)} = \frac{-5-(-1)}{-5-(5)} \]

OpenStudy (anonymous):

@DanJS which number do i plug in?

OpenStudy (anonymous):

This helps to find \((0,y_1)\), which I guess it the \(y\) intercept.

OpenStudy (danjs):

(x1,y1) = any of the 2 points given

OpenStudy (danjs):

y - (-1) = (2/5)(x - 5)

OpenStudy (danjs):

y + 1 = (2/5)x - 2

OpenStudy (danjs):

when x = 0 what is y?

OpenStudy (anonymous):

0? idk

OpenStudy (danjs):

For the intercepts, you set x=0 and find y, and set y=0 and find x (0,y) (x,0) in y+1 = (2/5)x - 2

OpenStudy (anonymous):

does this look correct? y+1=(2/5)(0)-2 y+1=(2/5)-2 y+1-1=(2/5)-2-1 y=2/5-1 am i on the right track or no?

OpenStudy (danjs):

When x = 0 you get y + 1 = (2/5)(0) -2 y = -3

OpenStudy (danjs):

2/5 times 0 is zero

OpenStudy (anonymous):

ok sorry missed that

OpenStudy (danjs):

So the point (0,-3) is the y intercept

OpenStudy (danjs):

Now set y = 0 and find x in y + 1 = (2/5)x - 2

OpenStudy (anonymous):

0+1=2/5(x)-2 1=2/5(x)-2 right track?

OpenStudy (danjs):

yes, now multiply everything by 5 to get rid of that fraction

OpenStudy (anonymous):

5=x-2

OpenStudy (danjs):

5 = 2x - 10

OpenStudy (anonymous):

oops just realised that

OpenStudy (danjs):

so x = 15 / 2

OpenStudy (anonymous):

ok so now it would be 5=2x-10 5+10=2x-10+10

OpenStudy (danjs):

yes, then divide by 2

OpenStudy (anonymous):

7.5

OpenStudy (danjs):

so the point (15/2 , 0) is the x intercept

OpenStudy (danjs):

yeah or 7.5

OpenStudy (anonymous):

ok thank you so much!!

OpenStudy (danjs):

summary: 1)find the equation for the line 2) set x = 0 and find y ; (0,y) 3) set y = 0 and find x ; (x,0)